# Repetition Improves Language Model Embeddings

### v1

@@ 4 Methodology
To finetune the model, we optimize the SimCSE loss with in-batch and mined hard negatives. Since this is standard, we defer discussion of this to Appendix D. Each batch is constructed by sampling a dataset from our set of training dataset, and then collecting examples from only this dataset. We use GradCache to train with a large batch size (2048) with limited GPU memory (Gao et al., 2021a). We train with LoRA instead of full finetuning, with $r=16$ and $\alpha=16$. We choose $\tau=1/50$ and a learning rate of $8\times 10^{-4}$. We use the Mistral-7B instruction-tuned model as a backbone (Jiang et al., 2023a). Our choices aim to be consistent with prior literature (Wang et al., 2023; Su et al., 2022; Zhang et al., 2023a).
4.3 Massive Text Embedding Benchmark
For evaluation, we use the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022). For this paper, we focus on the English-language subset of the benchmark. MTEB is composed of a collection of 56 datasets that are grouped into different embedding tasks: classification, clustering, pair classification, reranking, retrieval, sentence similarity (STS), and summarization. The goal is to construct general purpose embeddings that are useful for solving each of the tasks. More information about MTEB is specified in Appendix A.1.
For the fine-tuning setting, we evaluate on the entire English-language subset. In the zero-shot setting, for convenience, we only evaluate on a subset of MTEB. We describe this subset in Appendix D.
5 Experiments
Strategy Model Pool  Clas. P. Cls. Clus. Retr. STS Rera. Average
Main results:
Echo (ours) Mistral 7B Mean   64.06 75.26 27.02 23.61 72.40 60.00 55.07
Classical Mistral 7B Mean   58.21 73.87 23.85 20.35 56.97 54.44 45.88
Prior work:
Summarization Mistral 7B Last 66.01 81.82 26.48 19.13 70.13 66.24 54.96
Ablations:
Echo Mistral 7B Last   63.11 57.93 12.82 2.97 39.14 47.35 36.60
Classical Mistral 7B Last   58.23 46.64 13.51 2.60 33.97 46.51 32.52
Echo LLaMA 7B Mean   61.64 66.29 25.11 16.12 66.18 56.35 50.26
Classical LLaMA 7B Mean   56.61 68.46 23.22 18.63 56.49 53.26 44.65
Echo LLaMA 13B Mean   64.65 74.57 25.72 26.58 72.20 62.68 55.60
Classical LLaMA 13B Mean   58.50 65.06 24.22 18.92 57.47 56.38 45.15
Table 1: Zero-shot scores on MTEB tasks for Mistral-7B. We use a retrieval validation set (FiQA2018) to select the best prompt. Refer to Appendix C for the scores with alternative validation sets. Top: Comparison of echo embeddings to classical embeddings. Center: Summarization approach to constructing embeddings (Jiang et al., 2023b). Bottom: Ablations, including last-token pooling and LLaMA-2-{7B, 13B}.
5.1 Evaluation of Zero-shot Embeddings
We compare the performance of classical, echo, and summarization embeddings on MTEB tasks (Table 1). We validate using a retrieval dataset from MTEB (FiQA2018) as a validation set as described in Section 4.3. We report the scores using alternative validation sets in Appendix C.
Echo embeddings outperform classical embeddings zero-shot.
We see that echo embeddings outperform classical embeddings by a large margin: on average, by nearly 10 points for Mistral-7B. Further, this performance increase is consistent across every MTEB category, across models (LLaMA-2 vs Mistral), and across scale (7B vs 13B). This demonstrates that echo embeddings can significantly improve the performance of embeddings on real data, suggesting that the failure mode of classical embeddings that we describe in Section 3 can affect performance on real data.
Qualitative comparison of classical and echo embeddings.
In Section 3, we demonstrate that classical embeddings overestimate the similarity between examples which are superficially similar based on tokens that appear early in the sequence. To build intuition that this applies to realistic data, we present the sentence pair from STSBenchmark, a sentence similarity task from MTEB, in which echo embeddings reduce error the most:
[…]
For this example, notice, that the sentence pair $(x_{1},x_{2})$ on which echo embeddings improve error the most has exactly the property we identify as a failure mode for classical embeddings: the sentence is superficially similar for the first few tokens. On the other hand $(y_{1},y_{2})$ does not have this property.
Quantitative evaluation of the failure mode.
The above example builds intuition that, even on real data, classical embeddings fail to properly estimate similarity on examples which are superficially similar in the early tokens. We quantitatively measure the degree to which classical and echo embeddings fail on sentences which are similar for early tokens, and for sentences which are not. We find that classical embeddings systematically fail on examples which exhibit this structure, while echo embeddings do not. For convenience, we defer the discussion of these experiments and the results to Appendix C.1.
Last-token vs mean token pooling.
We find that last-token embeddings are substantially worse than mean token embeddings in the zero-shot setting, despite the fact that in principle, the last token in the sequence can encode information from all other tokens. In practice, it is clear that the last token does not encode sufficient information to achieve strong performance on MTEB in the zero shot setting.
Echo embeddings vs summarization.
We find that the average performance across the tested MTEB datasets is similar between echo and classical embeddings. Summarization does encourage the last token in the sequence to encode information about the entire sentence. We find that summarization is much more sensitive to the exact prompt while echo embeddings are robust to such minor variations (see Figure 5 in Appendix C). We suspect that echo embeddings are more robust as a result of more directly trying to encode bidirectional information into the embeddings.
5.2 Evaluation of Finetuned Embeddings
Strategy Model Pool   Clas. Clus. P. Cls. Rera. Retr. STS Average
Main results:
Echo (ours) Mistral 7B Last 77.43 46.32 87.34 58.14 55.52 82.56 64.68
Classical Mistral 7B Last  76.57 45.78 86.37 56.71 54.87 82.03 63.98
Prior work:
UAE-Large-V1 (MLM)  75.58 46.73 87.25 59.88 54.66 84.54 64.64
multilingual-e5-large (MLM)  77.56 47.10 86.19 58.58 52.47 84.78 64.41
bge-large-en-v1.5 (MLM)  75.97 46.08 87.12 60.03 54.29 83.11 64.23
udever-bloom-7b (autoregr.)  72.13 40.81 85.4 55.91 49.34 83.01 60.63
sgpt-5.8b (autoregr.)  68.13 40.34 82.00 56.56 50.25 78.10 58.93
e5-mistral-7b555e5-mistral-7b was recently released and leverages high quality synthetic data to achieve strong performance which is not publicly released. We report their performance, but we do not explicitly compare to them (Wang et al., 2023). (autoregr.)  78.47 50.26 88.34 60.21 56.89 84.63 66.63
Ablations:
Echo Mistral 7B Mean  77.00 44.94 87.73 58.30 55.11 82.52 64.22
Classical Mistral 7B Mean  76.26 42.68 86.31 57.58 53.75 81.53 62.96
Classical Mistral 7B-bidir. Last  76.70 45.94 88.15 57.23 54.96 82.42 64.23
Table 2: Finetuning scores on MTEB tasks. Top: Apples-to-apples comparison of echo embeddings and classical embeddings in which we use echo embeddings and classical embeddings with last-token pooling, with the same training setup. Center: Performance of recent open source embedding models, annotated by base model type, masked-language model or autoregressive. Bottom: Ablations for finetuning: using mean token embeddings (first two lines) and using a bidirectional architecture (last line).
Different embeddings on the MTEB leaderboard are often fine-tuned on different datasets. In order to perform an apples-to-apples comparison between embedding strategies, we fine-tune both echo and classical embeddings on the exact same datasets (described in Section 4.2). We report the results in Table 2. This table also includes a comparison to prior state-of-the-art methods using masked language models (MLM) and autoregressive language models. Further, we evaluate a number of ablations to determine the role of pooling strategy and architecture.
Echo embeddings outperform classical embeddings after finetuning.
We observe that echo embeddings consistently outperform classical embeddings on each category even after finetuning. Hence, the fundamental gap we find between classical and echo embeddings in Section 3 and in our zero-shot experiments persist after fine-tuning.
Comparison to prior state-of-the-art models.
We present comparisons to both prior MLM-based embeddings and prior autoregressive-language-model embeddings, listing the open-source models from the MTEB leaderboard. It is striking that MLMs vastly outperformed autoregressive models until recently. Our classical embeddings outperform the previous-best autoregressive language model. This is a result of using the strongest public 7B parameter language model (Mistral) and more fine-tuning data. However, despite these choices, classical embeddings do not outperform prior MLM-based approaches, perhaps because MLMs encode bidirectional context unlike classical embeddings from autoregressive models. Interestingly, echo embeddings allow us to close the gap to achieve state-of-the-art (on average) with an autoregressive model compared to prior open-sourced models on the leaderboard that used MLMs. A recent exception is the concurrent work by Wang et al. (2023) that use synthetic data to improve classical embeddings extracted from Mistral-7B. Their synthetic data is not publicly available, but the apples-to-apples comparison between classical and echo embeddings we performed suggests that echo embeddings could provide gains further gains over the numbers reported in (Wang et al., 2023) when fine-tuning with synthetic data.
Why doesn’t last-token pooling close the gap?
Since classical last-token embeddings can attend to every other token, they do not necessarily suffer from the failure mode that we highlighted in Section 3. Last token does not reliably capture relevant information in a zero-shot setting, but this could have been bridged via fine-tuning. It is thus surprising that, even after finetuning last-token embeddings that could (in principle) encode any embedding function, echo embeddings outperform classical embeddings. We identify two hypotheses that may explain this performance gap: (1) While last-token embeddings can attend to every token, the intermediate representations of earlier tokens cannot. If last-token pooling derives information from the internal representations of earlier tokens, by attending to these representations, last-token classical embeddings may still suffer from the failure mode of the earlier tokens. (2) If the post-finetuning performance benefits from the model initialization point, last-token classical embeddings may suffer: in Section 5.1 we show that last-token embeddings achieve poor zero-shot performance. We leave it to future work to explore these hypotheses. We do, however, observe that the gap between last- and mean token echo embeddings is smaller than the gap between last- and mean token classical embeddings, suggesting that echo embeddings can especially improve the quality of mean token embeddings.
Can we relax autoregressive language models to a bidirectional architecture and fine-tune?
To test the role of architecture, we finetune Mistral-7B on the same setup described in Section 4.2 but modified the architecture so as to remove the causal attention mask. While the initial weights are identical to Mistral-7B, this new model has bidirectional attention. We observe that the performance of bidirectional classical embeddings are better than our standard (causal) classical embeddings, but worse than echo embeddings. This suggests that the architecture alone is not sufficient to improve performance.
@@ 6 Related Work
Sentence embeddings.
Dense low-dimensional vectors representing textual semantics has been widely studied and applied. Early approaches involved computing embeddings for individual words (Hinton, 1984; Rumelhart et al., 1986; Elman, 1990; Mikolov et al., 2013; Pennington et al., 2014). Later work aims to compute dense vectors representing the semantics of entire sequences by combining or composing word vectors (Le and Mikolov, 2014; Iyyer et al., 2015; Kiros et al., 2015; Socher et al., 2011; Tai et al., 2015; Wang et al., 2016; Wieting et al., 2015). Khattab and Zaharia (2020) propose to use late interaction between document and query vectors to improve retrieval performance. Reimers and Gurevych (2019) propose S-BERT which takes a pretrained BERT (Devlin et al., 2018) and trains with a triplet loss on anchor sentences, semantically similar positive examples, and semantically dissimilar negative examples. More recent approaches typically adopt this approach with different pretrained models and a contrastive objective such as InfoNCE (Oord et al., 2018) or SimCSE (Gao et al., 2021b). Ni et al. (2021a) with Ni et al. (2021b) extend this approach to the T5 architecture (Raffel et al., 2020). Multiple papers use an additional unsupervised contrastive objective (Izacard et al., 2021; Wang et al., 2022). Other papers propose including prompts to improve task-specific embedding performance (Jiang et al., 2022; Su et al., 2022). Some work combines multiple of these training objectives and approaches (Xiao et al., 2023a; Li et al., 2023). Notably, except for the most recent approaches, nearly all embeddings were based upon bidirectional architectures that were often pretrained with a masked-language modeling objective.
Next-token language modeling for embeddings.
A series of papers aim to construct high quality embeddings from autoregressive large language models. Multiple papers apply the fine-tuning approach of S-BERT to language models but using a trained GPT (Radford et al., 2018) as the backbone architecture (Muennighoff, 2022; Zhang et al., 2023a). Ma et al. (2023) adopts this approach but for LLaMA-2 (Touvron et al., 2023). Jiang et al. (2023b) extracts embeddings by asking a language model to summarize the input sentence. Wang et al. (2023) is concurrent to our work and improves embeddings by adding synthetic training data and trains on Mistral (Jiang et al., 2023a).
Zero-shot embeddings.
Most recent sentence embeddings research has focused on improving finetuning. Reimers and Gurevych (2019) demonstrates that without finetuning, BERT has low-quality embeddings. To our knowledge, Jiang et al. (2023b) is the only paper that constructs zero-shot embeddings for autoregressive language models.
@@ 7 Conclusion
We have compared classical and echo embeddings in a toy example, on real data in the zero-shot setting, and after finetuning. With the toy data, we identified a failure mode of autoregressive classical embeddings, which we have shown can be recovered with echo embeddings. Our result motivates the development of higher quality embeddings which are important in retrieval applications.
In addition, until recently, masked language models largely dominated the MTEB leaderboard, despite often having an order of magnitude fewer parameters, having been trained on substantially less data, and performing worse on other benchmarks of interest to the natural language processing community. While our results do not explicitly explain the surprising success of masked language models, they do suggest that next-token language models suffer from an inherent drawback that may have stifled their performance until they became performant enough to compensate for this shortcoming. We believe that our embedding strategy achieves the best of both worlds: we gain the capability of next-token language models while recovering from the failure mode that next-token language models do not encode information about future tokens in their contextualized token embeddings.
8 Limitations
Despite the success of echo embeddings, the method has limitations. First, while echo embeddings achieve superior performance to classical embeddings, they require double the inference cost to pass two copies of the input sequence to the model. Though this is double the training cost for a fixed number of training steps, we show in Appendix D that echo embeddings achieve improved performance even when matching compute. Second, we do not fully explain why echo embeddings are improved in comparison to classical embeddings after finetuning even though there is no representational limitation. We leave it to future work to understand the exact underlying mechanisms for this improvement.
Acknowledgements
@@ A.2 Base Model HuggingFace IDs
Appendix B Echo Embeddings: Additional Information
@@
measures the similarity between the embeddings of $x$ and $y$ for any embedding function $\phi\colon\mathcal{X}\to R^{d}$. The cosine similarity is used for our experiments in Sections 3, and as the similarity function for training in 5. All MTEB datasets use cosine similarity to compute similarity with the exception of the classification datasets, in which similarity is not explicitly measured, and the clustering datasets, which use Euclidean distance,
@@
Throughout the paper, we use a variety of different prompts to construct embeddings. In Section C, we demonstrate that for zero-shot embeddings, the exact wording or template used as a prompting strategy does not have a strong effect on the performance of MTEB tasks, with the exception of for the summarization approach. This implies, in general, that classical embeddings and echo embeddings should be robust to the exact choice of prompts. The important component of echo embeddings is instead the structure: the input text should be repeated twice when computing embeddings, and the embeddings should be taken over the second occurrence of the input text.
@@
We provide a subset of the toy data from Section 2. For Structure 1, the data is given in Table 4. For Structure 2, the data is given in Table 5. For Structure 3, the data is given in Table 6. In all cases, the data is generated by GPT4. The data from Structure 1 is generated from the following GPT4 prompt, and the other structures are generated from minor variations on this:
Together, we need to generate sentence triplets. Each triplet will have the following form:
@@
- sentence 3 should be extremely similar to sentence 1 and semantically equivalent, but slightly re-worded
@@
"sentence1": "I like to eat apples and bananas but I really hate almost every other fruit.",
"sentence2": "I like to eat apples and bananas and I also enjoy also every other fruit",
"sentence3": "I like two fruits: apples and bananas but I hate nearly all fruits other than these.",
@@
Appendix C Additional Zero-shot Results
In this section, we describe the omitted methodology and results for the zero-shot section.
Prompt sampling procedure.
  1. 1.
    Choose an instruction. For classical embeddings, we choose from {Write, Say, Complete, Explain}. For echo embeddings, we choose from {Repeat, Rewrite, Rephrase, Fill in the blank}. For summarization, we choose from {Summarize, Categorize, Understand, Analyze}.
  2. 2.
    Choose a wording for the instruction. For example, if we chose “Say” as the instruction, then we would choose from {Say a sentence, Say a paragraph, Say something, Say a response, Say a query, Say a prompt}. For summarization, we also choose a second part of the wording, as the summarization strategy requires that the summary be in one word: {in one word, with a single word, succinctly with one word, in a unique one-word way, in a single word, in a word}.
  3. 3.
    Choose a separator, which include colons, commas, newlines.
  4. 4.
  5. 5.
    Classical prompts have the form: “{instruction} {separator} {prefix} SS”.
  6. 6.
[…]
  7. 7.
    Summarization prompts have the form: “{instruction0} {separator} {prefix} $S$ {instruction1} {separator}”.
For classical, we choose the prompts:
Write a sentence I] S
Write a prompt!
(I) S
Write some text
PROMPT-S
SAY A PARAGRAPH | SENTENCE 0] S
Say a query QUERY: S
Say a sentence!
[A] S
COMPLETE THE PROMPT Text (1) S
Complete the query SENTENCE 0) S
Complete the sentence:-S
Explain a query text 0 S
Explain a prompt | Sentence 1> S
EXPLAIN A SENTENCE Prompt (1) S
For echo, we choose the prompts:
Repeat The Paragraph.
query 1) S.
query 2) S
Repeat the response.
1) S.
AGAIN 2) S
REPEAT THE SENTENCE :: PROMPT
S :: RESPONSE
S
Rewrite the query | QUERY (A) S | (B) S
Rewrite the text. SENTENCE A) S. B) S
Rewrite the response | query A] S | query B] S
Rephrase the sentence:@S:Again@S
Rephrase The Sentence!
Text <> S!
Answer <> S
REPHRASE THE QUERY Sentence a) S Answer b) S
Fill in the blank in the prompt:
Query a) S:
Query b) S
FILL IN THE BLANK IN THE RESPONSE | Sentence A) S | Sentence B) S
Fill in the blank in the paragraph.
Text | S.
Response | S
For summarization, we use the prompts:
SUMMARIZE THE QUERY.
Prompt: SIN A WORD.
Summarize the sentence!
PROMPT <1> SSuccinctly With One Word!
SUMMARIZE THE PARAGRAPH. PROMPT (0) SIN A WORD.
CATEGORIZE THE PROMPT query
SWith a single word
Categorize the query | prompt [1] Sin a word |
CATEGORIZE THE SENTENCE.
Prompt <1> SIN A WORD.
Understand the sentence
@SIn a single word
Understand The Prompt:QUERY [0] Sin a single word:}
UNDERSTAND THE PARAGRAPH:Text I] SSuccinctly with one word:}
Analyze the sentence.
Sentence SIn A Unique One-word Way.
Analyze the response! query a> SIN A UNIQUE ONE-WORD WAY!
Analyze The Prompt
Sentence a> SIn a unique one-word way
Subset of MTEB for zero-shot evaluation.
We evalaute on the following subset of MTEB: FiQA2018, SCIDOCS, SciFact, NFCorpus, TwitterSemEval2015, TwitterURLCorpus, ImdbClassification, AmazonReviewsClassification, TweetSentimentExtractionClassification, MTOPDomainClassification, TwentyNewsgroupsClustering, BiorxivClusteringS2S, MedrxivClusteringS2S, StackOverflowDupQuestions, AskUbuntuDupQuestions, SciDocsRR, BIOSSES, STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark, and SICK-R.
We plot the sensitivity of classical, repetition, and summarization to different choices of prompts for different models in Figures 56, and 7. We also extend to plotting on all tested datasets individually in Figures 89, and 10. We observe that summarization is highly sensitive to the exact prompt used. However, neither classical nor echo were particularly sensitive in any case. Consistently, mean token pooling outperformed last token pooling by a large factor.
Evaluation of zero-shot results with different validation sets.
We include the zero results of validation using different MTEB datasets. For validation, we select one dataset from each category, as follows: classification: ImdbClassification; Pair Classification: TwitterSemEval2015; Clustering: TwentyNewsgroupsClustering; Retrieval: FiQA2018; STS: STSBenchmark, Reranking: StackOverflowDupQuestions. We plot these results for different models in Tables 89, and 10. We observe similar results across different validation sets, with minor variations in the performance. In addition, we the performance of each dataset when the prompts have been validated with FiQA2018 in Tables 1112, and 13.
C.1 Validating the connection between our synthetic data experiments and real data.
In Section 3, we hypothesized that classical embeddings would overestimate similarity on sentences where the first half of the sentence are similar, and that echo embeddings would recover from this failure mode. In order to test this hypothesis, we exact a set of examples from the STS datasets included in the MTEB benchmark in which the first half of the sentence is similar, and measure the degree to which the similarity is overestimated.
As a control, we also select points which are similar in the second half of the sentence, and measure the degree to which similarity is overestimated. By comparing the degree to which sentences which are similar in the first half are overestimated in similarity, and the degree to which sentences which are similar in the second half are overestimated, then we can identify if classical embeddings overestimate similarity in specifically sentences which are similar in the first half. Thus, under our hypothesis, we expect that, for classical embeddings, sentences which are similar in the first half are overestimated in similarity more than sentences that are similar in their second half. On the other hand, we expect that, for echo embeddings, the degree to which similarity is over- or underestimated is independent of whether the sentences are similar in the first or second half of the sentence.
Measuring sentence similarity estimation error.
Results.
Qualitative examples.
Appendix D Additional Finetuning Results
Training Datasets.
GPUs.
Training a model takes approximately two days on 4 A100 GPUs.
Instructions for finetuning datasets.
Models on MTEB leaderboard.
We compare our implementation of classical and echo embeddings to state-of-the-art approaches on MTEB. Namely, we display results for UAE-Large-V1 (Li and Li, 2023), multilingual-e5-large (Wang et al., 2024), bge-large-en-v1.5 (Xiao et al., 2023b), udever-bloom-7b (Zhang et al., 2023a), sgpt-5.8b (Muennighoff, 2022), e5-mistral-7b (concurrent work) (Wang et al., 2023).
Additional ablations.
We plot additional ablations, including ablating the role of instructions during training and evaluation, as well as providing an evaluation at step 280 (out of 720 total steps), which is approximately $1/3$ of the duration of training (Table 15). We note that echo embeddings still outperform classical embeddings in this setting.
Performance over training time.
Computational benefits of echo embeddings.
From Table 15, we observe that even after approximately $1/3$ of the total training duration (less than $1/2$), echo embeddings achieve performance higher than classical embeddings achieve after an entire epoch (Table 2). Echo embeddings requires twice the computational cost of classical embeddings. However, this result suggests that despite this additional cost per embedding, training with echo embeddings can save on training costs by requiring less than half an epoch of training to outperform classical embeddings. Further, since each data point is only seen once, it implies that echo embeddings are much more data efficient than classical embeddings, which may be helpful when data is costly or difficult to acquire.
All results.
NLI Retrieve semantically similar text
Squad Retrieve Wikipedia passages that answer the question
TriviaQA Retrieve Wikipedia passages that answer the question
Table 3: Instructions for finetuning datasets.
Training objective.
@@ Appendix D Additional Finetuning Results
What kind of insulation is it? What kind of floors are above? 0.0 "Tony’s not feeling well," Spurs coach Gregg Popovich said. We’re thrilled to be up 3-2,” Coach Gregg Popovich said Wednesday. 1.6

### v2

@@ 4.4 Evaluation of Fine-tuned Embeddings
Training datasets and optimization: We train on a collection of publicly available datasets that are standard training datasets in the embedding literature. We list and describe each of the datasets in Appendix F. To fine-tune the model, we optimize the SimCSE loss with in-batch and mined hard negatives. Since this is standard, we defer discussion of this to Appendix F. Each batch is constructed by sampling a dataset from our set of training dataset, and then collecting examples from only this dataset. We use GradCache to train with a large batch size (2048) with limited GPU memory (Gao et al., 2021a). We train with LoRA instead of full finetuning, with $r=16$ and $\alpha=16$. We choose $\tau=1/50$ and a learning rate of $8\times 10^{-4}$.
Compute-matching echo embeddings. Echo embeddings require twice the compute of classical embeddings. Similar to the zero-shot setting, we reduce the inference-time cost of echo embeddings by halving the length of the input. However, in the training setting, we additionally must match the train compute. We opt to train for half as many steps as classical embeddings, as this is the most straightforward way to match the compute. The compute-matched echo embeddings in Table 5 refer to the setting where we match compute for both training and inference.
Classical embeddings with bidirectional attention. In addition to our comparison to classical embeddings with causal attention, we train a model for which we replace the causal attention with bidirectional attention.
4.5 Fine-tuning Results
Echo embeddings yield high quality embeddings in the fine-tuning setting. We present the main fine-tuning results in Table 5. Similar to the zero-shot setting, compute-matched echo embeddings outperform classical embeddings with both architectures: causal and bidirectional. When additional compute is available, echo embeddings improve performance further. This suggests that the fundamental gap between classical and echo embeddings that we identified in Section 3 persists even after fine-tuning. We find consistent results when evaluating across model classes and scales (Table 7).
Figure 4: Zero-shot setting: MTEB-mini score of zero-shot echo embeddings and classical embeddings with bidirectional attention with fixed compute (token) budget. We use the zero-shot Mistral 7B from Table 1. The standard token budget for MTEB is 512.
Comparison to LLM2Vec. We compare to the supervised variant of LLM2Vec. The LLM2Vec method relaxes the attention of the base autoregressive model to bidirectional and requires two fine-tuning stages: (1) fine-tuning with an unsupervised masked-language-modeling objective (2) supervised fine-tuning for embeddings. Despite the complexity of the LLM2Vec method and additional fine-tuning, echo embeddings nonetheless approximately matches the performance with identical fine-tuning and inference-time compute (Table 7). Echo embeddings thus offers itself as a much simpler method that does not require modifying the architecture nor additional fine-tuning stages, with nearly identical performance.
Comparison to GritLM. We also compare to GritLM (Muennighoff et al., 2024) in Table 7. This approach, which simultaneously trains an embedder and a generative model, finds the bidirectional attention relaxation to be crucial. We find that compute-matched echo embeddings perform similarly the variant of GritLM that has been trained on similar data as ours (public data), even though echo embeddings use a causal architecture.
The role of pooling. We investigate the role of pooling in Appendix F. In the fine-tuning setting, the choice of pooling strategy has a much weaker effect than in the zero-shot setting. Consistent with prior work (Wang et al., 2023; Zhang et al., 2023a), we find that last-token pooling, with the addition of a trainable end-of-sentence token embedding, slightly outperforms mean pooling.
Figure 5: Fine-tuning setting: MTEB-mini score of fine-tuned echo embeddings and classical embeddings with bidirectional attention with fixed compute (token) budget. We use the trained Mistral-7B models from Table 5. The standard token budget for MTEB is 512.
4.6 Addressing the Inference-Time Cost of Echo Embeddings
One potential drawback of echo embeddings is the computational cost of the method at both training and inference time. Naively, echo embeddings requires approximately twice the compute of classical embeddings, as the input is repeated. We have so far seen that in an entirely compute-matched training and inference time setting, echo embeddings outperforms classical embeddings (Table 5). In this section, we expand this analysis to consider how echo embeddings perform under different inference-time compute budgets. We explore the role of training under different budgets in Appendix F.
4.7 The role of compute
The role of inference-time compute in the zero-shot setting. In Section 4.2, we demonstrated that by halving the number of input tokens that are encoded by echo embeddings, thus matching compute with the classical approach to emebddings, we achieve stronger performance than classical embeddings. Here, we investigate the role of compute in further detail by plotting the performance of echo embeddings as a function of the total number of tokens encoded (Figure 4). We find that when the budget is greater than approximately 64 tokens, echo embeddings outperforms Mistral-7B classical embeddings with bidirectional attention. For very small computational budgets (fewer than 64 tokens), we find that Mistral-7B classical embeddings with bidirectional outperform echo embeddings. Since the standard budget for MTEB is 512 tokens, echo embeddings likely offer stronger performance in most settings. Over all compute budgets, echo embeddings outperform causal attention classical embeddings.
The role of inference-time compute in the fine-tuning setting. The fine-tuning setting exhibits a similar trend: for very small computational budgets, echo embeddings perform worse than classical embeddings, but as the budget increases echo embeddings outperform classical embeddings (Figure 5). In the fine-tuning setting, echo embeddings outperform classical embeddings for budgets greater than 128 tokens, which is still much smaller than the standard MTEB budget (512 tokens).
@@ 5 Related Work
Sentence embeddings. Dense low-dimensional vectors representing textual semantics has been widely studied and applied. Early approaches involved computing embeddings for individual words (Hinton, 1984; Rumelhart et al., 1986; Elman, 1990; Mikolov et al., 2013; Pennington et al., 2014). Later work aims to compute dense vectors representing the semantics of entire sequences by combining or composing word vectors (Le & Mikolov, 2014; Iyyer et al., 2015; Kiros et al., 2015; Socher et al., 2011; Tai et al., 2015; Wang et al., 2016; Wieting et al., 2015). Khattab & Zaharia (2020) propose to use late interaction between document and query vectors to improve retrieval performance. Reimers & Gurevych (2019) propose S-BERT which takes a pretrained BERT (Devlin et al., 2018) and trains with a triplet loss. More recent approaches typically adopt this approach to different pretrained models (Raffel et al., 2020; Ni et al., 2021b; Raffel et al., 2020) and contrastive objectives (Oord et al., 2018; Gao et al., 2021b). Other papers propose including prompts to improve task-specific embedding performance (Jiang et al., 2022; Su et al., 2022). Some work combines multiple of these training objectives and approaches (Xiao et al., 2023a; Li et al., 2023). Most recent sentence embeddings research has focused on improving finetuning. Reimers & Gurevych (2019) demonstrates that without finetuning, BERT has low-quality embeddings. To our knowledge, Jiang et al. (2023b) is the only paper that constructs zero-shot embeddings for autoregressive language models.
Next-token language modeling for embeddings. A series of papers aim to construct high quality embeddings from autoregressive large language models. Most closely related, BehnamGhader et al. (2024) proposed LLM2Vec which adds an additional unsupervised training step on a masked-language-modeling objective.Multiple papers apply the fine-tuning approach of S-BERT to language models but using a trained GPT (Radford et al., 2018) as the backbone architecture (Muennighoff, 2022; Zhang et al., 2023a). Ma et al. (2023) adopts this approach but for LLaMA-2 (Touvron et al., 2023b). Jiang et al. (2023b) extracts embeddings by asking a language model to summarize the input sentence. Wang et al. (2023) improves embeddings by adding synthetic training data and trains on Mistral (Jiang et al., 2023a). Muennighoff et al. (2024) adds an additional objective to maintain generation performace, but requires bidirectional attention. Li & Li (2023b) modifies the architecture of a causal language model to be bidirectional.
Language modeling with bidirectional information. Xu et al. (2023) finds that repetition can improve the performance of autoregressive language models for reasoning tasks. arora2024just finds that repetition can be useful to improve recall for recurrent architectures.
@@ 6 Conclusion
We have demonstrated that the common assumption that bidirectional architectures are crucial for high quality embedding models is false. Our method, echo embeddings, prompts an autoregressive model to act as an autoecoder, which enables the model to encode bidirectional information. Our results highlight the importance of bidirectional information—we find that the inability to encode bidirectional information leads to the failure of classical-style embeddings in autoregressive models. Further, our method demonstrates that by carefully choosing the task that the language model is prompted to perform at inference time is a promising approach for improving embeddings. We hope that our work will inspire future research towards understanding how inference-time changes can impact embeddings.
Acknowledgments
@@ A.2 Base Model HuggingFace IDs
Appendix B Societal Considerations
In this section we discuss the societal impact of our work and the safeguards we have taken to mitigate negative impacts.
B.1 Impact
We intend for and believe that this work will have a positive societal impact. In particular, language models are known to suffer from limitations such as hallucinations and biases. In the context of retrieval-augmented generation, we believe that higher quality embeddings can lead to more accurate retrievals, which can in turn lead to more accurate generations. This can help mitigate the impact of hallucinations and biases in language models. We also believe that the improved embeddings can be used in a variety of downstream tasks, such as information retrieval, question answering, and summarization, which can have a positive impact on society. Nonetheless, we acknowledge that not all uses of our work will be necessarily positive. For example, our work can be used to improve the performance of language models in generating fake news or misinformation. Second, we release a training dataset that has been aggregated from a set of other open source datasets that might nonetheless contain harmful content. We believe that the positive impacts of our work outweigh the negative impacts, and we are committed to working towards ensuring that our work is used for positive purposes.
Appendix C Echo Embeddings: Additional Information
@@
measures the similarity between the embeddings of $x$ and $y$ for any embedding function $\phi\colon\mathcal{X}\to R^{d}$. The cosine similarity is used for our experiments in Sections 3, and as the similarity function for training in 4.4. All MTEB datasets use cosine similarity to compute similarity with the exception of the classification datasets, in which similarity is not explicitly measured, and the clustering datasets, which use Euclidean distance,
@@
Throughout the paper, we use a variety of different prompts to construct embeddings. In Section D, we demonstrate that for zero-shot embeddings, the exact wording or template used as a prompting strategy does not have a strong effect on the performance of MTEB tasks, with the exception of for the summarization approach. This implies, in general, that classical embeddings and echo embeddings should be robust to the exact choice of prompts. The important component of echo embeddings is instead the structure: the input text should be repeated twice when computing embeddings, and the embeddings should be taken over the second occurrence of the input text.
@@
We provide a subset of the toy data from Section 2. For Structure 1, the data is given in Table 8. For Structure 2, the data is given in Table 9. For Structure 3, the data is given in Table 10. In all cases, the data is generated by GPT-4. The data from Structure 1 is generated from the following GPT-4 prompt, and the other structures are generated from minor variations on this: Together, we need to generate sentence triplets. Each triplet will have the following form:
@@
- Sentence 3 should be extremely similar to sentence 1 and semantically equivalent, but slightly re-worded.
@@
”sentence1”: ”I like to eat apples and bananas but I really hate almost every other fruit.”,
”sentence2”: ”I like to eat apples and bananas and I also enjoy also every other fruit”,
”sentence3”: ”I like two fruits: apples and bananas but I hate nearly all fruits other than these.”,
@@
What kind of insulation is it? What kind of floors are above? 0.0 ”Tony’s not feeling well,” Spurs coach Gregg Popovich said. We’re thrilled to be up 3-2,” Coach Gregg Popovich said Wednesday. 1.6